4) How many grams of oxygen are present in 9.51 g of acetylsalicylic acid (CsHaO4), better known as aspirin?
Answer #1

molar mass of an aspirin molecule = 180g

mass of oxygen in an aspirin molecule =4*8g= 32g

so 180g of aspirin contains 32g of oxygen,so 1g of aspirin contains 32/180 g of oxygen

mass of oxygen in 9.51g of aspirin =9.51*(32/180) =1.69g

thank you